OpenStudy (anonymous):

what is x+3y=5 in polar form

OpenStudy (anonymous):

i think it is \[r \cos \theta-3rsin \theta=5\]

OpenStudy (jdoe0001):

ahemm hint \({\color{brown}{ x}}+3{\color{blue}{ y}}=5\qquad \qquad \begin{cases} {\color{brown}{ x}}=rcos(\theta)\\ {\color{blue}{ y}}=rsin(\theta) \end{cases}\) replace, and solve for "r" since polar equations are written as an "r" function in \(\theta\) terms

OpenStudy (jdoe0001):

yeap one sec

OpenStudy (jdoe0001):

\({\color{brown}{ x}}+3{\color{blue}{ y}}=5\qquad \qquad \begin{cases} {\color{brown}{ x}}=rcos(\theta)\\ {\color{blue}{ y}}=rsin(\theta) \end{cases} \\ \quad \\ \quad \\ rcos(\theta)+3rsin(\theta)=5\implies r[cos(\theta)+3sin(\theta)]=5 \\ \quad \\ r=\cfrac{5}{[cos(\theta)+3sin(\theta)]}\)

OpenStudy (jdoe0001):

hmm rats, missing an "r" there on the 2nd line anyway \(rcos(\theta)+3rsin(\theta)=5\implies r[cos(\theta)+3rsin(\theta)]=5 \\ \quad \\ r=\cfrac{5}{[cos(\theta)+3sin(\theta)]}\)

OpenStudy (anonymous):

so how would i do this problem \[y=3x^2 +1\] and change it into polar

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

try posting it anew.... while I see what I can find someone else may know it, and I'll check as well

OpenStudy (anonymous):

okay